# Diffusion Models in 3D Vision: A Survey

### v1

@@ IV-F Scene Generation
Scene generation focuses on constructing complete 3D environments, including objects, textures, and spatial layouts. Diffusion models play a key role by learning to synthesize cohesive scenes with realistic spatial relationships between objects, lighting conditions, and textures. This has implications for virtual world creation, architectural design, and immersive experiences in gaming.
Kim et al.[112] study how to automatically generate generative models for complex open-world 3D environments. The authors first train a scene autoencoder to express image and pose pairs as a neural field. They then use a Latent Auto-Encoder to further compress the neural field representation, mapping voxel grids to a set of latent representations. Finally, they train a hierarchical diffusion model on the latent representations to complete the scene generation process.
Wynn et al.[113] address the issue that NeRF can severely restrict scene geometry and color fields when only a limited number of input views are available, which may lead to unrealistic rendering results. The authors propose a method that utilizes a Denoising Diffusion Model (DDM) as a learned prior to regularize NeRF, using the DDM to learn prior knowledge about the scene’s geometry and color. This DDM is trained on RGBD image patches from the synthetic Hypersim dataset.
Chung et al.[114] presents LucidDreamer, which maintains and expands its world model through the repeated application of two operations: ”Dreaming” and ”Alignment.” The Dreaming process involves generating geometrically consistent images and elevating these images into 3D space. The Alignment algorithm harmoniously integrates the newly generated 3D point cloud segments into the existing 3D scene.
Fridman et al.[115] propose a novel framework that combines a pre-trained text-to-image generation model with a monocular depth prediction model, optimizing during testing to achieve online 3D-consistent scene video generation.
Wu et al.[116] introduce a method called BlockFusion, which utilizes latent 3D plane extrapolation techniques to compress 3D planes into a latent plane space using a Variational Autoencoder (VAE). In this space, a denoising diffusion process is performed to generate high-quality 3D scenes that can be infinitely scalable.
Gao et al.[117] present GraphDreamer, a framework that generates compositional 3D scenes from scene graphs. By leveraging scene graph information, it effectively addresses the issues of attribute confusion and guidance collapse in text-to-3D generation.
Ran et al.[23] introduce a new generative model called LiDAR Diffusion Models (LiDMs), which aims to enhance the fidelity of LiDAR scenes by capturing realism through the incorporation of geometric priors. This approach improves pattern authenticity, geometric authenticity, and object authenticity, enabling the generation of LiDAR scenes that are closer to real-world data.
Tang et al.[118] propose DiffuScene, which aims to generate diverse and realistic indoor 3D scenes. DiffuScene is based on a denoising diffusion model that represents the scene as a collection of unordered objects, where each object is characterized by attributes including position, size, orientation, semantics, and geometric features. It synthesizes a 3D indoor object collection by denoising a series of unordered object attributes.
Po et al.[119] explore how to control the generation of different parts of a scene using text prompts and bounding boxes while ensuring natural transitions between these parts. The authors propose Locally Conditioned Diffusion, which allows users to control specific areas of an image through text prompts and corresponding bounding boxes.
@@
Zhang et al.[120] present a framework called Text2NeRF, which combines text-to-image diffusion models with NeRF to generate 3D scenes with complex geometric structures and high-fidelity textures from text prompts. This method successfully enhances the realism and multi-view consistency of the generated scenes through progressive scene refinement and update strategies, as well as depth alignment strategies.
@@ IV-I Depth Estimation
Ji et al.[142] propose DDP, a conditional diffusion model framework designed for various dense visual prediction tasks. By effectively extending the denoising diffusion process within modern perception pipelines, DDP achieves competitive performance across multiple benchmarks without the need for task-specific designs or architectural customization. Notably, this method is one of the earliest to apply DDPMs to depth map prediction.
Li et al.[143] introduces DADP, a framework that combines diffusion models to improve depth estimation in sparsely annotated autonomous driving scenes. DADP includes a noise predictor and a depth predictor where the noise predictor enforce regional and holistic spatial structures by predicting noise components, while the depth predictor can be any single-image depth estimation model.
Zhang et al.[144] utilize Depth2Underwater ControlNet to generate a new underwater depth estimation dataset from terrestrial depth data using Stable Diffusion. This method combines the accuracy of terrestrial depth information with the realism of underwater scenes, and the dataset can be used to train underwater depth estimation models.
Patni et al.[22] propose a new monocular depth estimation model called ECoDepth, which uses a diffusion model as the backbone network. It leverages global image priors generated by a pretrained Vision Transformer (ViT) model as input to provide more detailed contextual information.
Saxena et al.[145] propose a model called DDVM, a diffusion model designed for monocular depth estimation. It combines self-supervised pretraining with supervised fine-tuning. DDVM is first pretrained on self-supervised image-to-image translation tasks, then undergoes supervised training on RGB-D data.
Ke et al.[146] introduce Marigold, a method based on stable diffusion models, which achieves state-of-the-art monocular depth estimation across multiple real-world datasets by fine-tuning on synthetic data without access to real depth samples. The authors only perform fine-tuning on the denoising U-Net, preserving the rich visual priors of the pre-trained model.

### v2

@@ IV-F Scene Generation
Scene generation focuses on constructing complete 3D environments, including objects, textures, and spatial layouts. Diffusion models play a key role by learning to synthesize cohesive scenes with realistic spatial relationships between objects, lighting conditions, and textures. This has implications for virtual world creation, architectural design, and immersive experiences in gaming. Scene generation can generally be divided into image-guided, text-guided, and scene graph-guided generation as shown in Figure 8.
Among different guidance, image-guided scene generation is the simplest and most intuitive. To automatically generate generative models for complex open-world 3D environments, Kim et al.[113] first train a scene autoencoder to express image and pose pairs as a neural field. They then use a Latent Auto-Encoder to further compress the neural field representation, mapping voxel grids to a set of latent representations. Finally, they train a hierarchical diffusion model on the latent representations to complete the scene generation process. Wynn et al.[114] address the issue that NeRF can severely restrict scene geometry and color fields when only a limited number of input views are available. The authors propose a method that utilizes a DDPM , which is trained on RGBD image patches from the synthetic Hypersim dataset [154], as a learned prior to regularize NeRF to learn prior knowledge about the scene’s geometry and color.. Chung et al.[115] presents LucidDreamer, which maintains and expands its world model through the repeated application of two operations: ”Dreaming” and ”Alignment.” The Dreaming process involves generating geometrically consistent images and elevating these images into 3D space. The Alignment algorithm harmoniously integrates the newly generated 3D point cloud segments into the existing 3D scene. Wu et al.[116] introduce a method called BlockFusion, which utilizes latent 3D plane extrapolation techniques to compress 3D planes into a latent plane space using a VAE. In this space, a denoising diffusion process is performed to generate high-quality 3D scenes that can be infinitely scalable. Ran et al.[23] introduce LiDAR Diffusion Models (LiDMs), which aims to enhance the fidelity of LiDAR scenes by capturing realism through the incorporation of geometric priors. This approach improves pattern authenticity, geometric authenticity, and object authenticity, enabling the generation of LiDAR scenes that are closer to real-world data.
@@
Compared to image-guided generation, text-guided scene generation offers greater flexibility and diversity. Fridman et al.[117] propose a novel framework that combines a pre-trained text-to-image generation model with a monocular depth prediction model, optimizing during testing to achieve online 3D-consistent scene video generation. Po et al.[118] explore how to control the generation of different parts of a scene using text prompts and bounding boxes while ensuring natural transitions between these parts. The authors propose Locally Conditioned Diffusion, which allows users to control specific areas of an image through text prompts and corresponding bounding boxes. Zhang et al.[119] present a framework called Text2NeRF, which combines text-to-image diffusion models with NeRF to generate 3D scenes with complex geometric structures and high-fidelity textures from text prompts. This method successfully enhances the realism and multi-view consistency of the generated scenes through progressive scene refinement and update strategies, as well as depth alignment strategies.
Compared to images and text, scene graphs [155] provide structured information that allows models to better understand and utilize the semantic relationships between objects. This effectively captures the multi-level relationships within a scene, resulting in generation outcomes that are more hierarchical and logical. Gao et al.[120] present GraphDreamer, a first-of-its-kind framework that generates compositional 3D scenes from scene graphs. By leveraging scene graph information, it effectively addresses the issues of attribute confusion and guidance collapse in text-to-3D generation. to generate diverse and realistic indoor 3D scenes, Tang et al.[121] propose DiffuScene, which is based on a denoising diffusion model that represents the scene as a collection of unordered objects, where each object is characterized by attributes including position, size, orientation, semantics, and geometric features. It synthesizes a 3D indoor object collection by denoising a series of unordered object attributes.
Zhai et al.[122] propose EchoScene, which employs a dual-branch diffusion model. Each branch contains multiple denoising processes, with an equal number of nodes at each point in the scene graph. The model exchanges information at each denoising step through an information echo scheme, combining these updates using graph convolution. This scheme ensures that the denoising process is influenced by a global understanding of the scene graph, thereby facilitating the generation of globally consistent scenes.
@@ IV-I Depth Estimation
Current research in depth estimation faces a shortage of training data, prompting some researchers to attempt using DDPMs to generate realistic depth map samples. Zhang et al.[147] utilize Depth2Underwater ControlNet to generate a new underwater depth estimation dataset from terrestrial depth data using Stable Diffusion. This method combines the accuracy of terrestrial depth information with the realism of underwater scenes, and the dataset can be used to train underwater depth estimation models. To address the challenges in monocular depth estimation tasks caused by complex and adverse factors, Tosi et al.[148] generates user-defined new scenes with a series of challenges and corresponding depth information, using the latest text-to-image diffusion models with depth-aware control. Moreover, Ke et al.[151] introduce Marigold, a method based on stable diffusion models, which achieves state-of-the-art monocular depth estimation across multiple real-world datasets by fine-tuning on synthetic data without access to real depth samples. The authors only perform fine-tuning on the denoising U-Net, preserving the rich visual priors of the pre-trained model.
Pre-training techniques are also extensively utilized to assist DDPMs in depth estimation, helping to overcome the shortage of training data. Patni et al.[22] leverage global image priors generated by a pretrained Vision Transformer (ViT) [156] model as input to provide more detailed contextual information. Saxena et al.[150] propose a model called DDVM, which combines self-supervised pretraining with supervised fine-tuning. DDVM is first pretrained on self-supervised image-to-image translation tasks, then undergoes supervised training on RGB-D data.
To address the issue of researchers relying on synthetic datasets, Shim et al.[149] integrate diffusion models into a domain-adaptive depth estimation framework, extracting domain-invariant structural content. It utilizes depth-consistent style transfer to estimate attention maps, resulting in improved performance in both indoor and outdoor environments compared to existing methods. As one of the earliest to apply DDPMs to depth map prediction, Ji et al.[144] using multi-task techniques, propose a conditional diffusion model framework designed for various dense visual prediction tasks. By effectively extending the denoising diffusion process within modern perception pipelines, DDP achieves competitive performance across multiple benchmarks without the need for task-specific designs or architectural customization. Li et al.[145] introduces DADP, a framework that combines diffusion models to improve depth estimation in sparsely annotated autonomous driving scenes. DADP includes a noise predictor and a depth predictor where the noise predictor enforce regional and holistic spatial structures by predicting noise components, while the depth predictor can be any single-image depth estimation model. Instead of directly diffusing the ground-truth (GT) depth, Duan et al.[146] introduce a new approach called DiffusionDepth, which learns to reverse the diffusion of its own refined depth back to a random depth distribution, which overcomes the challenge of applying generative models in sparse GT depth scenarios.